OpenStudy (anonymous):

simplify sin^2θ/cosθ+cosθ

geerky42 (geerky42):

\[\dfrac{\sin^2 \theta}{\cos\theta} + \cos \theta = \dfrac{\sin\theta}{\cos \theta} \sin\theta + \cos \theta = \tan \theta \sin \theta + \cos\theta\]

geerky42 (geerky42):

Hmm. Better plan: \[\dfrac{\sin^2 \theta}{\cos\theta} + \cos\theta = \dfrac{\sin^2 \theta}{\cos\theta} + \dfrac{\cos^2 \theta}{\cos \theta} = \dfrac{\sin^2 \theta + \cos^2 \theta}{\cos\theta} = \dfrac{1}{\cos\theta} \\ \Huge = \boxed{\sec\theta}\]

geerky42 (geerky42):

Is this clear? @Zeecoop